OpenStudy (anonymous):

Rewrite as a logarithmic equation. 2^4 = (a-b)

OpenStudy (anonymous):

4log2=log(a-b)

OpenStudy (solomonzelman):

I meant \[a^b=c---> Log_ac=b\]

OpenStudy (anonymous):

OR Log2 base (A-B) = 4

OpenStudy (anonymous):

Awesome sauce :) lol

OpenStudy (anonymous):

please give medal

OpenStudy (solomonzelman):

@mkuma36 fishing for medals like this is against the code.

OpenStudy (anonymous):

how ?

OpenStudy (solomonzelman):

How is it against the code, ask any moderator or ambassador. If you want to give someone a medal, click best response next to the answer.